OpenStudy (anonymous):

The distributive property allows you to say that 3(x-1) = 3x- ? A. 1 B. -1 C. 3 D. -3

OpenStudy (sidsiddhartha):

distributive property means \[a(b+c)=ab+ac\] so in ur problem\[3(x-1)=3x-3\] so what do u think is the correct option?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

or a?

OpenStudy (sidsiddhartha):

how can it be 1